HYDRATED HCl COMPLEXES, (HCl)_{m}(H_{2}O)_{n}, IN HELIUM NANODROPLETS: OH STRETCHING MODES

SEUNG JUN LEE and \underline{MYONG YONG CHOI}, Department of Chemistry (BK21) and Research Institute of Natural Science, Gyeongsang National University, Jinju, 660-701, Korea; DIM-ITRY SKVORTSOV and ANDREY F. VILESOV, Department of Chemistry, University of Southern California, Los Angeles, CA 90089, USA.

Infrared laser spectroscopy has been used to characterize the solvation of HCl with small number of water molecules in helium nanodroplets. Complexes of (HCl)_{m}(H_{2}O)_{n} with m = 1 – 3 and n = 1 – 3 have been identified in the free OH stretching and bonded OH spectral range of 3000 – 3900 cm^{−1}. The assignment of the infrared vibrational bands of the clusters is aided by pick-up pressure dependence experiments and ab initio calculations.